The area of a triangle is 6 square feet. The base is 3 times the height. What is the base of the triangle?.
Answer by Boreal(15235)

Area=1/2 bh
6=1/2bh
12=bh
height is x, base 3x
12=3x^2
x^2=4; x=2, positive root only.
base is 6 feet.
height is 2 feet
area is 1/2 ^*2=6 sq ft as check.